OpenStudy (anonymous):

Six friends are at the school carnival. They all want to go on the same ride, but only three people can go on the ride at one time. How many different combinations of three riders are possible? A.18 B.9 C.20 D.40

OpenStudy (anonymous):

i think its A. but i would get a second opinion

OpenStudy (anonymous):

you are being asked for "6 choose 3" which is \[\frac{6\times 5\times 4}{3\times 2}=20\]

OpenStudy (anonymous):

hmm im split between 2 answers who should i go with?

OpenStudy (anonymous):

@satellite73 r u 100% sure?

OpenStudy (anonymous):

you should learn how to compute "n choose k" for yourself it is written as \(\binom{n}{k}\) or sometimes \(_nC_k\)